OpenStudy (anonymous):

solve (1/9)= 27^(3x-5)

OpenStudy (anonymous):

-2/3 = 3x-5 5-2/3 = 3x = 4 1/3 or 13/3 x = 13/3 * 1/3 = 13/9.

OpenStudy (campbell_st):

\[\frac{1}{9} = 3^{-2}\] and \[27^(3x-5)= 3^{3\times(3x - 5)}\] equate the powers and solve for x -2 = 9x - 15 13 = 9x x = 13/9